OpenStudy (anonymous):

Find the magnitude ||v|| and direction angle θ, to the nearest tenth of a degree, for the given vector v. v = -4i - 3j B. 7; 216.9° C. 5; 216.9° D. 5; 36.9°

OpenStudy (anonymous):

@MaimiGirl help with this last one?

OpenStudy (anonymous):

do you know what the definition of the magnitude is?

OpenStudy (anonymous):

sorry i am blank on this one

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I'll sit back and learn from @RBauer4

OpenStudy (anonymous):

me too :) cause Idk about this

OpenStudy (anonymous):

We can define the magnitude as the square root of the sum of the components squared. For example, our vector is this caseis\[\left(\begin{matrix}-4 \\ -3\end{matrix}\right)\]

OpenStudy (anonymous):

So the magnitude would be \[\sqrt{ (-4)^2 + (-3)^2}\]

OpenStudy (anonymous):

ok what would I do next

OpenStudy (anonymous):

To find the azimuthal angle of this vector, we can analyze its scalr product against i. There is an equation, <x , y> = ||x||*||y||cos(theta), where theta is the angle between vectors x and y.

OpenStudy (anonymous):

Let x equal our vector and let y be i.

OpenStudy (anonymous):

Ok, i am with you

OpenStudy (anonymous):

So let's solve for theta.

OpenStudy (anonymous):

ok how do you do that i am sorry i don't know how

OpenStudy (anonymous):

<x,y> = -4*1 + -3*0 = -4 , ||x|| = 5, ||y|| = 1 so cos(theta) = -4/5

OpenStudy (anonymous):

ok i am starting to understand

OpenStudy (anonymous):

What was A for this question?

OpenStudy (anonymous):

I got it wrong so i didn't post it

OpenStudy (anonymous):

but this was A. 5; 233.1°

OpenStudy (anonymous):

i see. I ask because our vector points in the third quadrant. So our angle will be greater than 180, but less than 270.

OpenStudy (anonymous):

I was thinking C. but i wasn't really sure

OpenStudy (anonymous):

C doesn't make since calculations wise. If we were to measure from 180 to theta, then we would get about 36.9.

OpenStudy (anonymous):

so it would be 5; 36.9°

OpenStudy (anonymous):

I think you might be right I would go for C, since 360 - 143.1 = 216.9

OpenStudy (anonymous):

That was what I was thinking but i wasn't 100% sure